# Multifractal critical phase driven by coupling quasiperiodic systems to electromagnetic cavities

### v1

@@ IV Conclusions
We have analyzed the phase transition using both the average and minimum inverse participation ratios, $\overline{\text{IPR}}$ and $\text{IPR}_{\text{min}}$, respectively. While the former determines when a fraction of the states cease to be extended, the latter quantifies the presence of at least one extended state. We have established that, for weak coupling with the electromagnetic cavity, the critical behavior is similar to that for the standard AAH model, with all eigenstates undergoing a phase transition at the same critical point. However, at intermediate or strong coupling, this single critical point broadens into an intermediate region, where $\text{IPR}_{\text{min}}$ vanishes, but $\overline{\text{IPR}}$ remains finite. Further, the size of this intermediate region increases with larger couplings and is strongly dependent on the photon frequency. We have also proposed an energy-resolved phase diagram, which displays the critical points at different energy ranges. As the main result, we observed that the effects of coupling with the cavity are more pronounced at higher energies, with the states at the bottom of the spectrum being hardly affected.
The nature of the intermediate region was examined through a multifractal analysis, by means of the exponent $\alpha_{min}$, which provides a measure of how a given wavefunction spatially spreads through the medium. For the standard AAH model, it shows that the eigenstates are fractal only at the critical point. A similar analysis for the field-coupled case provides evidence of fractal states within the entire intermediate region. We have also investigated the critical properties of the eigenstates by the probability distributions of $\alpha_{min}$ and the IPR. Both distributions are quite broad, strongly suggesting a mixture of extended, localized, and critical states within the intermediate region. That is, the coupling with the electromagnetic field drives the single critical point of the standard AAH model into a critical region. These findings unveil the impact of the coupling of optical cavities on critical transport phenomena and paves the way for further investigations of the influence of cavity-coupling in other metal-insulator transitions. Finally, we hope that the present results may find applications in the design of materials with specific critical transport properties, and in the study of the physics of disordered and quasicrystals structures Krishna Kumar et al. (2018); Xu et al. (2010); Chen et al. (2024).
@@ ACKNOWLEDGMENTS
We are grateful to C. Lewenkopf and L. Martin-Moreno for valuable discussions. J.F. thanks FAPERJ, Grant No. SEI-260003/019642/2022; RRdS acknowledges grants from CNPq [314611/2023-1] and FAPERJ [E-26/210.974/2024 - SEI-260003/006389/2024]; N.C.C. acknowledges support from FAPERJ Grant No. E-26/200.258/2023 - SEI-260003/000623/2023, and CNPq Grant No. 313065/2021-7; T.F.M. acknowledges support from FAPERJ Grant No. E-26/202.518/2024 - SEI-260003/007717/2024. Financial support from the Brazilian Agencies Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq), Coordenação de Aperfeiçoamento de Pessoal de Ensino Superior (CAPES), and Instituto Nacional de Ciência e Tecnologia de Informação Quântica (INCT-IQ) is also gratefully acknowledged. F.A.P acknowledges support from CAPES, CNPq, and FAPERJ.

### v2

@@ IV Conclusions
We have analyzed the phase transition using both the average, typical, and minimum inverse participation ratios, $\overline{\text{IPR}}$, $\text{IPR}_{\rm typ}$, and $\text{IPR}_{\text{min}}$, respectively. While the former (average and typical) determine when a fraction of the states cease to be extended, the latter quantifies the presence of at least one extended state. We have established that, for weak coupling with the electromagnetic cavity, the critical behavior is similar to that for the standard AAH model, with all eigenstates undergoing a phase transition at the same critical point. However, at intermediate or strong coupling, this single critical point broadens into an intermediate region, where $\text{IPR}_{\text{min}}$ vanishes, but $\text{IPR}_{\rm typ}$ remains finite. Further, the size of this intermediate region increases with larger couplings and is strongly dependent on the photon frequency. We have also proposed an energy-resolved phase diagram, which displays the critical points at different energy ranges. As the main result, we observed that the effects of coupling with the cavity are more pronounced at higher energies, with the states at the bottom of the spectrum being hardly affected.
The nature of the intermediate region was examined through a multifractal analysis, by means of the exponent $\alpha_{min}$, which provides a measure of how a given wavefunction spatially spreads through the medium. For the standard AAH model, it shows that the eigenstates are fractal only at the critical point. A similar analysis for the field-coupled case provides evidence of fractal states within the entire intermediate region. We have also investigated the critical properties of the eigenstates by the probability distributions of $\alpha_{min}$ and the IPR. Both distributions are quite broad, strongly suggesting a mixture of extended, localized, and critical states within the intermediate region. That is, the coupling with the electromagnetic field drives the single critical point of the standard AAH model into a critical region.
The emergence of multifractal states induced by light-matter interactions remains a subject under debate. For instance, Ref. Botzung et al. (2020) reports the presence of multifractal states in three-dimensional disordered arrays of quantum dots under light-induced effects. Similarly, Ref. Dubail et al. (2022) examines the manifestation of such states in a Bethe lattice. However, this picture changes significantly when more realistic conditions are considered. As discussed in Ref. Allard and Weick (2022), when dealing with a dispersive cavity and multiple cavity modes – which are essential ingredients for accurately modeling polaritonic systems –, such multifractal states disappear. A comparable suppression is observed in interacting systems, as discussed in Ref. Mattiotti et al. (2024), in which any perturbation of the integrability of the model eliminates multifractal behavior. Therefore, the occurrence of critical states in the AAH model driven by light-matter interactions is an important step towards the understanding of metal-insulator transitions in the presence of optical cavities. Finally, we hope that the present results may find applications in the design of materials with specific critical transport properties, and in the study of the physics of disordered and quasicrystals structures Krishna Kumar et al. (2018); Xu et al. (2010); Chen et al. (2024).
@@ ACKNOWLEDGMENTS
We are grateful to C. Lewenkopf, L. Martin-Moreno, T. Allard, G. Weick and C. Genet for valuable discussions. J.F. thanks FAPERJ, Grant No. SEI-260003/019642/2022; RRdS acknowledges grants from CNPq [314611/2023-1] and FAPERJ [E-26/210.974/2024 - SEI-260003/006389/2024]; N.C.C. acknowledges support from FAPERJ Grant No. E-26/200.258/2023 - SEI-260003/000623/2023, CNPq Grant No. 313065/2021-7, and Serrapilheira Institute Grant No. R-2502-52037; T.F.M. acknowledges support from FAPERJ Grant No. E-26/202.518/2024 - SEI-260003/007717/2024. F.A.P acknowledges support from CAPES, CNPq, and FAPERJ. Financial support from the Brazilian Agencies Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq), Coordenação de Aperfeiçoamento de Pessoal de Ensino Superior (CAPES), and Instituto Nacional de Ciência e Tecnologia de Informação Quântica (INCT-IQ) is also gratefully acknowledged.